OpenStudy (anonymous):

what is x when 116+4x+32=180

OpenStudy (anonymous):

4x = 180-116-32 = 32 x = 8

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

it says that the right answer was 64

OpenStudy (anonymous):

second time it says the right answer is wrong. something is out of whack

OpenStudy (anonymous):

what does the question say exactly?

OpenStudy (anonymous):

are you trying to find the area of something?

OpenStudy (anonymous):

pls recheck yr question

OpenStudy (radar):

64 is the 4x+32.

OpenStudy (anonymous):

this ismy next queston. Given parallelogram LBRW and line BR=15 and WL=3x-9, what is the value of x?

OpenStudy (anonymous):

guess you have to know something else yes?

OpenStudy (anonymous):

lb = wl so x = 8

OpenStudy (anonymous):

thats all i got

OpenStudy (anonymous):

br = wl meant to say

OpenStudy (anonymous):

Next. Given paralleglogram ABCD BA=32 and CD=9x+5, what is the value of x?

OpenStudy (anonymous):

solve 9x +5 = 32 and you\ve got it

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you soooo much

OpenStudy (anonymous):

yw